OpenStudy (surana):

Find the maximum or minimum of the following quadratic function: y=x^2 -16

OpenStudy (jdoe0001):

so.... any idesa where the vertex of that one is at?

OpenStudy (surana):

\[x^{2}\] It's supposed to look like that.

OpenStudy (surana):

I think the vertex is X.

OpenStudy (jdoe0001):

well... have you covered quadratics yet?

OpenStudy (surana):

Yes, but it's not my strong point.

OpenStudy (jdoe0001):

well let's take a peek .... \(\large { y=(x-{\color{brown}{ h}})^2+{\color{blue}{ k}}\\ x=(y-{\color{blue}{ k}})^2+{\color{brown}{ h}}\qquad\qquad vertex\ ({\color{brown}{ h}},{\color{blue}{ k}}) }\) notice that, notice where the vertex is at

OpenStudy (jdoe0001):

\(\bf y=x^2-16\implies y-{\color{blue}{ 0}}=(x-{\color{brown}{ 0}})^2-16\) notice the vertext on that one and the -16 is just a transformation of \(x^2\) -16 means, from that vertex, you move down 16 units

OpenStudy (surana):

It's at K?

OpenStudy (jdoe0001):

well... actually. hmmm hold the mayo... I'm mixing up two forms there...

OpenStudy (jdoe0001):

yes, just use the "k" form the vertex form

OpenStudy (jdoe0001):

lemme rewrite that quick

OpenStudy (jdoe0001):

\(\bf y=x^2-16\implies y=(x-{\color{brown}{ 0}})^2{\color{blue}{ -16}}\) there, now you can see where the vertex is at :)

OpenStudy (surana):

So the vertex is sixteen?

OpenStudy (jdoe0001):

the \(x^2\ has a positive coefficient meaning the parabola goes upwards a maximum point means a "hump" a minimum point, means a "burrow" in this case the parabola has a "burrow" not a "hump" thus it has a minumum then